y''-2y'-8y=4e2x-21e-3x

1
Expert's answer
2021-08-08T18:00:14-0400

Solution;

Auxiliary equation is ;

"m^2-2m-8=0"

"m^2-4m+2m-8=0"

"m(m-4)+2(m-4)=0"

"(m+2)(m-4)=0"

m=-2,4

"y_h=C_1e^{4x}+C_2e^{-2x}"

The particular solution;

"y_1=e^{4x}"

"y_2=e^{-2x}"

"g(x)=4e^{2x}-21e^{-3x}"

"W(y_1,y_2)=det\begin{vmatrix}
 e^{4x}& e^{-2x} \\
 4e^{4x} & -2e^{-2x}
\end{vmatrix}" "=-2e^{2x}-4e^{2x}=-6e^{2x}"

"W_1=det\begin{vmatrix}
 0 & e^{-2x}\\
 4e^{2x}-21e^{-3x} & -2e^{-2x}
\end{vmatrix}=-e^{-2x}(4e^{2x}-21e^{-3x})=-1+21e^{-5x}"

"W_2=det\begin{vmatrix}
 e^{4x} & 0 \\
 4e^{4x} & 4e^{2x}-21e^{-3x}
\end{vmatrix}=e^{4x}(4e^{2x}-21e^{-3x})=4e^{6x}-21e^{x}"

Let;

"u_1=\int\frac{W_1}{W}dx=\int\frac{-1+21e^{-5x}}{-6e^{2x}}dx=\frac16\int e^{-2x}dx-\frac{21}6\int e^{-7x}dx"

"u_1=\frac{-e^{-2x}}{12}+\frac{e^{-7x}}{2}"

"u_2=\int\frac{W_2}{W}dx=\int\frac{4e^{6x}-21e^x}{-6e^{2x}}=\frac{-4}6\int e^{4x}dx+\frac72\int e^{-x}dx"

"u_2=-\frac16e^{4x}-\frac72e^{-x}"

Let;

"y_p=u_1y_1+u_2y_2"

"y_p=(\frac{e^{-7x}}{2}-\frac{e^{-2x}}{12})e^{4x}+(-\frac16e^{4x}-\frac72e^{-x})e^{-2x}"

"y_p=-\frac{e^{2x}}{4}-3e^{-3x}"

The general solution is;

"y=y_h+y_p"

"y=C_1e^{4x}+C_2e^{-2x}-\frac{e^{2x}}{4}-3e^{-3x}"